$SPAD/src/input images7.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
-- Color gallery page 7
)r conformal
-- Conformal maps
-- The map z +-> z + 1/z on the complex plane
-- The coordinate grid for the complex plane
f z == z
conformalDraw(f, -2..2, -2..2, 9, 9, "cartesian")
f z == z + 1/z
conformalDraw(f, -2..2, -2..2, 9, 9, "cartesian")
-- The map z +-> -(z+1)/(z-1)
-- This function maps the unit disk to the right half-plane, as shown
-- on the Riemann sphere.
-- The unit disk
f z == z
riemannConformalDraw(f, 0.1..0.99, 0..2*%pi, 7, 11, "polar")
-- The right half-plane
f z == -(z+1)/(z-1)
riemannConformalDraw(f, 0.1..0.99, 0..2*%pi, 7, 11, "polar")
-- Visualization of the mapping from the complex plane to the Riemann Sphere.
riemannSphereDraw(-4..4, -4..4, 7, 7, "cartesian")

———-

References^{[1] nothing}

4